# Free field construction of Heterotic string compactified on Calabi-Yau manifolds of Berglund-Hubsch type in the Batyrev-Borisov combinatorial approach

### v6

@@ 4 Berglund-Hubsch type CY manifolds and N=2 SCFT models
In constructing the vertex algebra of the left and right CY-subsectors of Heterotic string, following the work of BOR1 ; BOR2 , we use vertices corresponding to the points of the dual lattices $M$ and $N$. In fact, this is only true for vertices that correspond to bosonic states (in the left sector, they correspond to NS states). To construct fermion states we need to use in both sectors, left and right, vertices corresponding to the following extension of the $M$ and $N$ lattices, namely, $M\pm\frac{1}{2}\vec{a^{+}}$ and $N\pm\frac{1}{2}\vec{a^{-}}$.
@@
Following to Borisov BOR1 ; BOR2 we define the Vertex algebra of Heterotic string as the cohomology with respect to the diffential $D_{\vec{1},\vec{1}}$ defined as the sum of the differntials $D_{\vec{m}}$ and $D_{\vec{n}}$, where
$D_{\vec{m}}=\oint du\ \vec{m}\cdot\vec{\Psi}^{-}(u)\ \exp(\vec{m}\cdot\vec{X^{-}}(u))=\sum_{i=1}^{5}m_{i}\oint du\exp(-iH_{i}+\vec{m}\cdot\vec{X^{-}}(u)),$ (14)
$D_{\vec{n}}=\oint du\ \vec{n}\cdot\vec{\Psi}^{+}(u)\ \exp(\vec{n}\cdot\vec{X^{+}}(u))=(\sum_{i=1}^{5}n_{i}\oint du\exp(iH_{i}+\vec{n}\cdot\vec{X^{+}}(u)),$
$D_{\vec{m}}^{2}=D_{\vec{n}}^{2}=\{D_{\vec{m}},D_{\vec{m}^{\prime}}\}=\{D_{\vec{n}},D_{\vec{n}^{\prime}}\}=\{D_{\vec{m}},D_{\vec{n}}\}=0;$
where $\vec{m}\in\Delta^{+}$ and $\vec{n}\in\Delta^{-}$. It was stated in BOR1 ; BOR2 that the elements $T$, $G^{\pm}$ and $J$ belong to the kernel of the operators $D_{\vec{m}}$ and $D_{\vec{n}}$.
@@ 9.4 The singlet supermultiplets of $E(8)\times E(6)$
$V_{\vec{\lambda},\vec{m},\vec{n}}(z,\bar{z})=\exp\left(-\phi+\imath{\vec{\lambda}}\cdot\vec{\tilde{H}}+\frac{1}{2}H^{L}_{CY}\right)\times\exp(\vec{m}\cdot\vec{X}^{-}+\vec{n}\cdot\vec{X}^{+}),$ (70)
@@
$V_{\vec{\lambda},\vec{m}}(z,\bar{z})=\exp\left(-\phi+\imath{\vec{\lambda}}\cdot\vec{\tilde{H}}+\frac{1}{2}H^{L}_{CY}\right)\times\sum_{i}m_{i}\exp{(-i\bar{H}_{i}+\vec{m}\cdot\vec{X}^{-})},$ (71)
@@
$V_{\vec{\lambda},\vec{n}}(z,\bar{z})=\exp\left(-\phi+\imath{\vec{\lambda}}\cdot\vec{\tilde{H}}+\frac{1}{2}H^{L}_{CY}\right)\times\sum_{i}n_{i}\exp{(i\bar{H}_{i}+\vec{n}\cdot\vec{X}^{+})},$ (72)

### v8

@@ 4 Berglund-Hubsch type CY manifolds and N=2 SCFT models
In constructing the vertex algebra of the left and right CY-subsectors of Heterotic string, following the work of Borisov BOR1 ; BOR2 , we use vertices corresponding to the points of the dual lattices $M$ and $N$. In fact, this is only true for vertices that correspond to bosonic states (in the left sector, they correspond to NS states). To construct fermion states we need to use in both sectors, left and right, vertices corresponding to the following extension of the $M$ and $N$ lattices, namely, $M\pm\frac{1}{2}\vec{a^{+}}$ and $N\pm\frac{1}{2}\vec{a^{-}}$.
@@
Namely, following BOR1 ; BOR2 , we define the Vertex algebra of Heterotic string as cohomology with respect to the sum of differentials $D_{\vec{u}_{i}}$ and $D_{\vec{v}_{j}}$
$D=\sum_{i=1}^{5}D_{\vec{u}_{i}}+\sum_{j=1}^{5}D_{\vec{v}_{j}},$ (14)
where
$D_{\vec{u}_{i}}=\oint dz\ \vec{u}_{i}\cdot\vec{\Psi}^{-}(z)\ \exp(\vec{u}_{i}\cdot\vec{X^{-}}(z)),$ (15)
$D_{\vec{v}_{j}}=\oint dz\ \vec{v}_{j}\cdot\vec{\Psi}^{+}(z)\ \exp(\vec{v}_{j}\cdot\vec{X^{+}}(z)).$
It was stated in BOR1 ; BOR2 that the elements $T$, $G^{\pm}$ and $J$ belong to the kernel of the operator $D$.
@@ 9.4 The singlet supermultiplets of $E(8)\times E(6)$
$V_{\vec{\lambda},\vec{m},\vec{n}}(z,\bar{z})=\exp\left(-\phi+\imath{\vec{\lambda}}\cdot\vec{\tilde{H}}\right)\times\exp(\vec{m}\cdot\vec{X}^{-}+\vec{n}\cdot\vec{X}^{+}),$ (71)
@@
$V_{\vec{\lambda},\vec{m}}(z,\bar{z})=\exp\left(-\phi+\imath{\vec{\lambda}}\cdot\vec{\tilde{H}}\right)\times\sum_{i}m_{i}\exp{(-i\bar{H}_{i}+\vec{m}\cdot\vec{X}^{-})},$ (72)
@@
$V_{\vec{\lambda},\vec{n}}(z,\bar{z})=\exp\left(-\phi+\imath{\vec{\lambda}}\cdot\vec{\tilde{H}}\right)\times\sum_{i}n_{i}\exp{(i\bar{H}_{i}+\vec{n}\cdot\vec{X}^{+})},$ (73)